# Hierarchical Position Embedding of Graphs with Landmarks and Clustering for Link Prediction

### v1

@@ 8. Related Work
Link Prediction. Earlier methods for link prediction used heuristics (Adamic and Adar, 2003; Zhou et al., 2009) based on manually designed formulas. GNNs were subsequently applied to the task, e.g., GAE (Kipf and Welling, 2016) is a graph auto-encoder which reconstructs adjacency matrices combined with GNNs, but cannot distinguish isomorphic nodes. SEAL (Zhang and Chen, 2018) is proposed as structural link representation by extracting enclosing subgraphs and learning structural patterns of those subgraphs. The authors demonstrated that higher-order heuristics can be approximately represented by lower-order enclosing subgraphs thanks to $\gamma$-decaying heuristic. LGLP (Cai et al., 2021) used the graph transformation prior to GNN layers for link prediction, and Walk Pooling (Pan et al., 2022) proposed to learn subgraph structure based on random walks. However, the aforementioned methods need to extract enclosing subgraphs of edges and compute their node embeddings on the fly. CFLP (Zhao et al., 2022) is a counterfactual learning framework for link prediction to learn causal relationships between nodes. However, its time complexity is $O(N^{2})$ for finding counterfactual links with nearest neighbors.
Distance- and Position-based Methods. P-GNN (You et al., 2019) proposed position-aware GNN based on distances for injecting positional information into node embeddings. P-GNN focuses on realizing Bourgain’s embedding (Bourgain, 1985) guided by Linial’s method (Linial et al., 1995) and performs message computation and aggregation based on distances to random subset of nodes. By contrast, we judiciously select representative nodes in combination with graph clustering and use the associated distances. Laplacian positional encodings (Belkin and Niyogi, 2003; Dwivedi et al., 2020) use eigenvectors of graph Laplacian as positional embeddings in which positional features of nearby nodes are encoded to be similar to one another. Graph transformer was combined with positional encoding learned from Laplacian spectrum  (Dwivedi and Bresson, 2021; Kreuzer et al., 2021). However, transformers with full attention have high computational complexity and do not scale well for link predictions in large graphs. Distance encoding (DE) as node labels was proposed and its expressive power was analyzed in (Li et al., 2020). In (Zhang et al., 2021), the authors analyzed the effects of various node labeling tricks using distances. However, these two methods do not utilize distances as positional information.
@@ Appendix A Proofs
We claim that $p^{\lambda}_{ij}(s)$ can be expressed as (14). The first summation in the bracket of (14) counts all possible walks of length $s$ from $i$ to $j$ and sums up their probabilities. The second summation in the bracket of (14) counts all possible walks from $i$ to $j$ but never visits $\lambda$. Thus the subtraction in the bracket counts all the walks from $i$ to $j$ visiting $\lambda$ at least once. Thus, expression (14) is the probability of the existence of walks of length $s$ from $i$ to $j$ via $\lambda$ from Lemma 1, e.g., $A_{k}$ in Lemma 1 corresponds to an event of a walk. The expression is asymptotically accurate in $N$, i.e., although Lemma 1 requires events $A_{k}$ be independent, and the same edge may participate between different $A_{k}$’s and induce correlation, the fraction of such correlations becomes negligible when $s\ll N$, as argued in (Fronczak et al., 2004) [^1] Alternatively, a simple proof can be done by using a counting argument: the ratio of the number of paths not visiting the same node more than once (which are uncorrelated paths) out of all the possible paths of length $s$ from node $i$ to $j$ is close to 1 if $s\ll N$..
@@
where the error term associated with exponential integral vanishes because $a/\log b$ is small, and $\gamma\approx 0.5772$ is the Euler’s constant. By applying the approximation and plugging in $a$ and $b$ to (24), we get expression (25) for the second term of (19).
@@ C.1. Graph Clustering Algorithm
C.2. Number of Clusters
@@
C.3. Node Centrality
Centrality DDI PubMed Cora Citeseer Facebook
Dataset vanilla GCN HPLC
CITATION2 282856 (ms) 352670 (ms)
COLLAB 1719 (ms) 2736 (ms)
DDI 3246 (ms) 3312 (ms)
PubMed 8859 (ms) 8961 (ms)
Cora 982 (ms) 1022 (ms)
Facebook 1107 (ms) 1243 (ms)
Dataset vanilla GCN HPLC
CITATION2 29562 (MB) 36344 (MB)
COLLAB 4988 (MB) 7788 (MB)
DDI 2512 (MB) 2630 (MB)
PubMed 2166 (MB) 3036 (MB)
Cora 1458 (MB) 6160 (MB)
Facebook 2274 (MB) 2796 (MB)
Appendix D Actual training time and GPU memory usage
Table 9 and 9 show the comparison and actual training time and GPU memory usage between vanilla GCN with HPLC. Note that HPLC already contains vanilla GCN as a component. Thus, one should attend to the additional resource usage incurred by HPLC in Table 9 and 9. Experiments are conducted with NVIDIA A100 with 40GB memory. We observe that, the additional space and time complexity incurred by HPLC relative to vanilla GCN is quite reasonable, demonstrating the scalability of our framework.
Appendix E Hyperparameters
Table 10. Dataset statistics.

### v2

@@ 8. Related Work
Link Prediction. Earlier methods for link prediction used heuristics (Adamic and Adar, 2003; Page et al., 1998; Zhou et al., 2009) based on manually designed formulas. GNNs were subsequently applied to the task, e.g., GAE (Kipf and Welling, 2016) is a graph auto-encoder which reconstructs adjacency matrices combined with GNNs, but cannot distinguish isomorphic nodes. SEAL (Zhang and Chen, 2018) is proposed as structural link representation by extracting enclosing subgraphs and learning structural patterns of those subgraphs. The authors demonstrated that higher-order heuristics can be approximately represented by lower-order enclosing subgraphs thanks to $\gamma$-decaying heuristic. LGLP (Cai et al., 2021) used the graph transformation prior to GNN layers for link prediction, and Walk Pooling (Pan et al., 2022) proposed to learn subgraph structure based on random walks. However, the aforementioned methods need to extract enclosing subgraphs of edges and compute their node embeddings on the fly. CFLP (Zhao et al., 2022) is a counterfactual learning framework for link prediction to learn causal relationships between nodes. However, its time complexity is $O(N^{2})$ for finding counterfactual links with nearest neighbors.
Distance- and Position-based Methods. P-GNN (You et al., 2019) proposed position-aware GNN to inject positional information based on distances into node embeddings. P-GNN focuses on realizing Bourgain’s embedding (Bourgain, 1985) guided by Linial’s method (Linial et al., 1995) and performs message computation and aggregation based on distances to random subset of nodes. By contrast, we judiciously select representative nodes in combination with graph clustering and use the associated distances. Laplacian positional encodings (Belkin and Niyogi, 2003; Dwivedi et al., 2020) use eigenvectors of graph Laplacian as positional embeddings in which positional features of nearby nodes are encoded to be similar to one another. Graph transformer was combined with positional encoding learned from Laplacian spectrum  (Dwivedi and Bresson, 2021; Kreuzer et al., 2021). However, transformers with full attention have high computational complexity and do not scale well for link predictions in large graphs. Distance encoding (DE) as node labels was proposed and its expressive power was analyzed in (Li et al., 2020). In (Zhang et al., 2021), the authors analyzed the effects of various node labeling tricks using distances. However, these two methods do not utilize distances as positional information.
@@ Appendix A Proofs
We claim that $p^{\lambda}_{ij}(s)$ can be expressed as
(11) $p^{\lambda}_{ij}(s)=1-\exp\left[-\left\{\sum_{v_{1}=1}^{N}\cdots\sum_{v_{s-1}=1}^{N}q_{iv_{1}}q_{v_{1}v_{2}}\cdots q_{v_{s-1}j}-\sum_{\begin{subarray}{c}v_{1}=1\\ v_{1}\neq\lambda\end{subarray}}^{N}\cdots\sum_{\begin{subarray}{c}v_{s-1}=1\\ v_{s-1}\neq\lambda\end{subarray}}^{N}q_{iv_{1}}q_{v_{1}v_{2}}\cdots q_{v_{s-1}j}\right\}\right]$
The first summation in the bracket of (11) counts all possible walks of length $s$ from $i$ to $j$ and sums up their probabilities. The second summation in the bracket of (11) counts all possible walks from $i$ to $j$ but never visits $\lambda$. Thus the subtraction in the bracket counts all the walks from $i$ to $j$ visiting $\lambda$ at least once. Thus, expression (11) is the probability of the existence of walks of length $s$ from $i$ to $j$ via $\lambda$ from Lemma 1, e.g., $A_{k}$ in Lemma 1 corresponds to an event of a walk. The expression is asymptotically accurate in $N$, i.e., although Lemma 1 requires events $A_{k}$ be independent, and the same edge may participate between different $A_{k}$’s and induce correlation, the fraction of such correlations becomes negligible when $s\ll N$, as argued in (Fronczak et al., 2004) [^1] Alternatively, a simple proof can be done by using a counting argument: the ratio of the number of paths not visiting the same node more than once (which are uncorrelated paths) out of all the possible paths of length $s$ from node $i$ to $j$ is close to 1 if $s\ll N$..
@@
where the error term associated with exponential integral vanishes because $a/\log b$ is small, and $\gamma\approx 0.5772$ is the Euler’s constant. By applying the approximation and plugging in $a$ and $b$ to (21), we get expression
for the second term of (16).
@@ C.1. Graph Clustering Algorithm
C.2. Node Centrality
Centrality DDI PubMed Cora Citeseer Facebook
Appendix D Actual training time and GPU memory usage
Table 8 and 8 show the comparison and actual training time and GPU memory usage between vanilla GCN with HPLC. Note that HPLC already contains vanilla GCN as a component. Thus, one should attend to the additional resource usage incurred by HPLC in Table 8 and 8. Experiments are conducted with NVIDIA A100 with 40GB memory. We observe that, the additional space and time complexity incurred by HPLC relative to vanilla GCN is quite reasonable, demonstrating the scalability of our framework.
Dataset vanilla GCN HPLC
CITATION2 282856 (ms) 352670 (ms)
COLLAB 1719 (ms) 2736 (ms)
DDI 3246 (ms) 3312 (ms)
PubMed 8859 (ms) 8961 (ms)
Cora 982 (ms) 1022 (ms)
Facebook 1107 (ms) 1243 (ms)
Dataset vanilla GCN HPLC
CITATION2 29562 (MB) 36344 (MB)
COLLAB 4988 (MB) 7788 (MB)
DDI 2512 (MB) 2630 (MB)
PubMed 2166 (MB) 3036 (MB)
Cora 1458 (MB) 6160 (MB)
Facebook 2274 (MB) 2796 (MB)
Appendix E Dataset
Dataset statistics is shown in Table 9.
Table 9. Dataset statistics.
Appendix F Hyperparameters
F.1. Number of Clusters
@@
F.2. Miscellaneous
Miscellaneous hyperparameters are shown in Table 11.